### v1

# Advancing Multi-Party Dialogue Systems with Speaker-ware Contrastive Learning

Abstract
Dialogue response generation has made significant progress, but most research has focused on dyadic dialogue. In contrast, multi-party dialogues involve more participants, each potentially discussing different topics, making the task more complex. Current methods often rely on graph neural networks to model dialogue context, which helps capture the structural dynamics of multi-party conversations. However, these methods are heavily dependent on intricate graph structures and dataset annotations, and they often overlook the distinct speaking styles of participants. To address these challenges, we propose CMR, a Contrastive learning-based Multi-party dialogue Response generation model. CMR uses self-supervised contrastive learning to better distinguish "who says what." Additionally, by comparing speakers within the same conversation, the model captures differences in speaking styles and thematic transitions. To the best of our knowledge, this is the first approach to apply contrastive learning in multi-party dialogue generation. Experimental results show that CMR significantly outperforms state-of-the-art models in multi-party dialogue response tasks.
Advancing Multi-Party Dialogue Systems with Speaker-ware Contrastive Learning
Zhongtian Hu1, Yiwen Cui1, Ronghan Li2, Meng Zhao3, Lifang Wang1, 1School of Computer Science and Engineering, Northwestern Polytechnical University, 2School of Computer Science and Technology, Xidian University, 3School of Artificial Intelligence and Big Data, Henan University of Technology ahxchzt@mail.nwpu.edu.cn, wanglf@nwpu.edu.cn
1 Introduction
In recent years, dialogue response generation has made notable progress Ide and Kawahara (2021); Liu et al. (2021c), largely due to advancements in language models. These improvements have enhanced the generation of coherent and contextually relevant responses in applications such as customer service, virtual assistants, and conversational agents Hu et al. (2024); Wu et al. (2022); Komeili et al. (2022). However, most research has focused on dyadic dialogues Cai and Cai (2022); Sun et al. (2021); Chen et al. (2022), involving only two participants. In contrast, multi-party dialogues, common in real-world scenarios like group discussions, meetings, and social media, present unique structural challenges. As shown in Figure 1, while dyadic dialogues follow natural speaker rotations, multi-party dialogues require careful management of information flow between multiple participants Gu et al. (2022b); Ganesh et al. (2023), posing significant challenges for response generation models.
Figure 1: This figure illustrates a multi-party dialogue involving four participants: John, Mike, Sarah, and Emma. Each participant engages in discussions on different topics, demonstrating the dynamics and complexity of multi-party conversations.
To address these complexities, many approaches utilize graph neural networks (GNNs) to model dialogue context Hu et al. (2019); Wang et al. (2020); Shi and Huang (2019). For example, Gu et al. (2022a) proposed the HeterMPC model, which uses utterance and interlocutor nodes with six meta relations to construct a heterogeneous graph that captures dialogue interactions. Similarly, Gu et al. (2023) introduced a fully connected graph to model utterance-interlocutor relationships, incorporating latent addressee inference. While GNNs effectively model the dependencies in multi-party dialogues, they have drawbacks. The complexity of graph structures and the need for extensive annotations, such as addressee labels Li and Zhao (2023), make these methods computationally expensive and challenging to scale in real-world applications.
Given these challenges, we propose CMR (Contrastive Learning-based Multi-Party Response Generation). Our approach uses contrastive learning to understand the structure and dynamics of multi-party dialogues without relying on complex graph structures or extensive annotations. The model follows a two-stage training process. In the first stage, only the encoder is trained. Through contrastive learning, the model contrasts utterances from different speakers within the same conversation and those from random speakers across the dataset. This helps the model distinguish between different facts and speaking styles, enhancing speaker awareness and leading to more relevant responses. In the second stage, the model is jointly trained on a response generation task with contrastive learning. Here, the model contrasts the dialogue context with generated responses (via beam search), the gold response, the target speaker’s previous utterance, and other speakers’ responses within the same conversation. This stage improves response quality by enhancing the model’s understanding of conversation themes.
We validated our approach using the Friends dataset. Both automatic evaluation results and human evaluation results demonstrate that our model significantly outperforms the baseline models.
Our primary contributions are as follows:
  • We propose CMR, a novel multi-party dialogue response generation model that leverages contrastive learning to better understand the structure and dynamics of multi-party dialogues. To the best of our knowledge, this is the first work on multi-party dialogue response generation based on contrastive learning.
  • We introduce a two-stage training process: the first stage focuses on distinguishing the speaking styles and facts among different participants through contrastive learning, and the second stage jointly trains the model with the response generation task and contrastive learning to improve the quality of responses and the model’s understanding of dialogue themes.
  • Both automatic evaluation results and human evaluation results on the Friends dataset demonstrate that our model significantly outperforms baseline models and achieves SOTA.
2 Related Work
Research on multi-party dialogue response generation is relatively sparse, but there are some pioneering works in this field. Additionally, we review some methods that combine contrastive learning with generation tasks.
2.1 Graph Neural Networks for Multi-Party Dialogue
Several studies have explored the use of GNNs to capture the intricate relationships in multi-party dialogues. For instance, GSN (Graph-Structured Network) introduced by Hu et al. (2019). models multi-party dialogues as graphs rather than sequences. GSN employs an utterance-level graph-structured encoder to model information flow along the graph and incorporates speaker information to improve performance Similarly, Gu et al. (2022a) proposed the HeterMPC model, which employs heterogeneous graph networks to model interactions between interlocutors and utterances using multiple types of nodes and edges. Furthermore, another approach by Gu et al. (2023) constructed fully connected graphs to model relationships between all utterances and interlocutors in a conversation. While addressee labels are not entirely unnecessary, their model is designed to infer latent addressees through a graph prediction model integrated into a Variational Inference framework, making it robust even when some addressee information is missing.
2.2 Enhanced Speaker-Aware Approaches
Ma et al. (2023) proposed an Enhanced Speaker-Aware Multi-Party Multi-Turn Dialogue Comprehension model that integrates speaker-related features at both the token and discourse levels. By applying a speaker mask for attention and using heterogeneous graph networks for comprehensive speaker-aware discourse clues, this model achieves state-of-the-art performance on datasets such as MolweniLi et al. (2020) and FriendsQAYang and Choi (2019). Liu et al. (2021b) proposed an Utterance-Aware and Speaker-Aware Multi-turn Dialogue Representation model (MDFN) that integrates utterance-level and speaker-level features using a mask attention mechanism. The MDFN achieve amazing performance on several Datasets for response selection task. These methods highlights the importance of incorporating speaker-aware signals to improve dialogue comprehension and response generation.
2.3 Contrastive Learning for Text Generation
Contrastive learning has been effectively applied to text generation tasks to enhance the quality of generated responses. For example, the work by An et al. (2022) on contrastive neural text generation demonstrates the benefits of using contrastive objectives to distinguish between similar and dissimilar contexts, thereby improving the coherence and relevance of generated text. Additionally, Liu et al. (2021a) introduced a topic-aware contrastive learning approach for abstractive dialogue summarization. By contrasting different dialogue contexts and responses, their model better captures thematic shifts and important content, leading to more accurate summaries.
3 Methodology
3.1 Task Formalization
Formally, let $D=\{({u_{1}},{s_{1}}),...,({u_{n}},{s_{m}})\}$ represent a multi-party conversation, which contains $n$ utterances and $m$ speakers, and $u_{i}$ is the utterance made by speaker $s_{j}$. Our objective is to develop a model $f$ that can generate a response $r$ for a specific speaker $s_{t}$ based on the context of the dialogue $D$. This can be expressed as: $r=f(D,s_{t})$. In the subsequent sections, we will describe our proposed model, CMR (Contrastive Learning-based Multi-Party Response Generation), which leverages contrastive learning in a two-stage training process to address these components effectively.
3.2 Model Overview
(a) The figure displays the training process of the stage one.
(b) The figure displays the training process of the stage two.
Figure 2: The figure displays the training processes of the CMR model. Stage One (a) shows the encoder training with contrastive learning, where the encoder processes the query and distinguishes between positive and negative samples. Stage Two (b) illustrates the joint training with response generation and contrastive learning, involving both the encoder and decoder. The combined loss of cross-entropy and InfoNCE enhances the model’s ability to generate contextually relevant responses and understand the dialogue context.
CMR is designed to effectively handle the complexities of multi-party dialogue response generation. The model employs a two-stage training process. In the first stage, the encoder is trained with contrastive learning to distinguish between the speaking styles and contexts of different participants. In the second stage, the model undergoes joint training with the response generation task and contrastive learning, enhancing its ability to generate contextually relevant and accurate responses.
3.3 Stage I: Enhancing Encoder Discrimination
The primary focus of stage I is to improve the encoder’s capacity to understand and differentiate between diverse dialogue contexts. Specifically, the encoder undergoes training using contrastive learning by contrasting the utterances of different speakers within the same conversation and those of random speakers from the entire dataset. Thus, the encoder can better distinguish between various speaking styles and related facts, enhancing its ability to capture the unique characteristics of each speaker.
3.3.1 Selection of Positive and Negative Samples
First, we randomly select an utterance from the conversation as the query $q$. The positive sample $k^{+}$ for this query is another utterance spoken by the same speaker within the same conversation. The negative samples ${k^{-}}$ consist of two parts: utterances spoken by other speakers within the same conversation, and utterances spoken by other speakers from the entire dataset, i.e., in-batch samples. It is worth noting that when selecting the query and positive and negative samples, we removed meaningless and dull replies, such as "yes yes yes," to ensure the quality and informativeness of the training data. The detailed methodology for data cleaning is provided in Appendix. By effectively distinguishing between these positive and negative samples, the encoder learns to better capture the nuances of multi-party dialogues.
3.3.2 Training Process of Stage I
The training objective for the first stage is to maximize the similarity between the query and its positive sample while minimizing the similarity between the query and its negative samples. The approach can help the model better understand "who says what" and the stylistic differences between different speakers. We achieve this by using the InfoNCE loss Oord et al. (2018) function.
$\mathcal{L}_{\text{stage-1 }}=-log\frac{exp(\frac{h_{q}\cdot h_{k^{+}}}{\tau})}{{\textstyle\sum_{i=1}^{N}}exp((\frac{h_{q}\cdot h_{k_{i}}}{\tau})+{\textstyle\sum_{j=1}^{T}}exp(\frac{h_{q}\cdot h_{k^{-}_{j}}}{\tau})}$ (1)
$h_{q}=Pooling(Encoder(q))$ (2)
$h_{k^{+}}=Pooling(Encoder(k^{+}))$ (3)
$h_{k^{-}}=Pooling(Encoder(k^{-}))$ (4)
where $h_{q}$ is the representation of the query, $h_{k^{+}}$ is the representation of the positive sample, $h_{k^{-}_{j}}$ represents the negative samples from the same conversation, $h_{k_{i}}$ represents the samples from-batch, and $\tau$ represents temperature. Pooling refers to utilizing the hidden state of the $<$/s$>$ token to represent the encoding of the entire sequence. Because our early experiments revealed that during generation, the decoder’s cross-attention in the T5 model focuses predominantly on the $<$/s$>$ token. Consequently, the $<$/s$>$ token effectively acts as an aggregator of context information, making it a suitable representation of the entire sequence. The parameters $T$ and $N$ denote the number of negative samples from the same conversation and the batch size, respectively.
3.4 Stage II: Joint Training with Response Generation and Contrastive Learning
After the encoder has been trained to distinguish between different speakers and their styles in Stage I, the next step involves joint training with response generation and contrastive learning, which aims to further refine the model’s ability to generate contextually relevant and coherent responses.
3.4.1 Selection of Positive and Negative Samples
As with traditional contrastive learning in generation tasks, we use the dialogue context as the query and the gold response as the positive sample. However, our approach introduces an innovative method for selecting negative samples, which consists of four parts:
  • Previous Utterance by the Target Speaker: The first part of the negative samples is the utterance immediately preceding the gold response spoken by the target speaker. In general, the utterance is highly related to the gold response but contextually different enough due to the progression of the dialogue. We believe the utterance serving as a negative sample can help the model understand the dialogue context by distinguishing between closely related but semantically distinct utterances.
  • Utterances by Other Speakers in the Current Conversation: The second part includes utterances from other speakers within the same conversation with query. These samples help the model grasp the context by understanding topic shifts and interactions between different participants.
  • From Batch Samples: The third part consists of negative samples drawn from the entire dataset (batch samples). These provide a broader context and ensure that the model generalizes well across different dialogues.
  • Model-Generated Beam Search Samples: The fourth part includes the model’s own beam search samples. Following An et al. (2022), these samples, generated by the model itself, are similar in structure to the gold response but differ in content. Including these helps the model refine its understanding and improve the quality of generated responses.
3.4.2 Training Process of Stage II
In Stage Two, the model is trained to generate responses while simultaneously leveraging contrastive learning to enhance the quality of these responses. For the response generation task, we use maximum likelihood estimation with cross-entropy loss. For the contrastive learning part, we use the InfoNCE loss, similar to Stage One. Thus, the training objective in Stage Two combines the response generation loss and the contrastive learning loss. The combined loss function can be expressed as:
$\mathcal{L}_{\text{stage-2 }}=\mathcal{L}_{\text{gen }}+\lambda\mathcal{L}_{\text{cl}}$ (5)
where $\mathcal{L}_{\text{gen }}$ is the response generation loss, calculated using cross-entropy loss between the generated response and the gold response. $\mathcal{L}_{cl}$ is the contrastive learning loss, and $\lambda$ is a hyperparameter that balances the two loss components.
The cross-entropy loss for response generation is defined as:
$\mathcal{L}_{\text{gen }}=-\sum_{t=1}^{T}\log P\left(r^{t}\mid r^{<t},D,s_{t}\right)$ (6)
Here, $r^{t}$ represents the token at position $t$ in the gold response, $r^{<t}$ represents the sequence of tokens in the gold response before position $t$. $D$ is the dialogue context, and $s_{t}$ is the target speaker. The cross-entropy loss measures the negative log-likelihood of the gold response given the dialogue context and the target speaker.
The contrastive learning loss $\mathcal{L}_{cl}$ is defined as:
$\mathcal{L}_{cl}=-log\frac{exp(\frac{h_{q}\cdot h_{k^{+}}}{\tau})}{exp(\frac{h_{q}\cdot h_{k^{+}}}{\tau})+{\textstyle\sum_{j=1}^{P+T+N+B}}exp(\frac{h_{q}\cdot h_{k^{-}_{j}}}{\tau})}$ (7)
where $h_{q}$ is the representation of the query, $h_{k^{+}}$ is the representation of the positive sample, and $h_{k^{-}_{j}}$ represents the negative samples. The representations $h_{q}$ and $h_{k}$ for the query and the samples are obtained in a manner similar to Stage One. The parameters $P,T,N,B$ denote the number of negative samples from four different sources.
The entire training process can be referenced in Appendix 1.
4 Experiments
4.1 Dataset
For evaluating the performance of our proposed CMR model, we considered two widely-used benchmarks in multi-party dialogue research: the Friends datasetShmueli and Ku (2019) and the Ubuntu IRC benchmarkLowe et al. (2015). The Friends dataset, which contains transcripts from the TV show "Friends," features interactions between six main characters with distinct and consistent personalities, providing a rich context for modeling speaking styles. These well-defined personas allow the model to generate dialogue that aligns with each character’s unique traits. In contrast, the Ubuntu IRC benchmark includes conversations from a large number of speakers, many of whom only contribute a single utterance. This makes it challenging to model individual speaking styles. Additionally, the Ubuntu dataset includes addressee tags and other labels that are not typically available in natural multi-party dialogues and differ from the objectives of our model. Given these factors, we determined that the Friends dataset is more suitable for evaluating our model’s ability. After data pre-processing, there are 5000 conversations for training and 653 conversations for testing.
4.2 Comparison Models
Models Metrics F1 BLEU-1 BLEU-2 BLEU-3 Rouge-L
BartLewis et al. (2020) 4.80 4.78 1.64 0.78 4.46
T5Raffel et al. (2020) 4.92 5.19 1.65 0.67 4.65
HeterMPCGu et al. (2022a) 5.13 5.30 1.41 0.40 4.74
MADNetGu et al. (2023) 6.17 6.00 1.54 0.63 5.95
CONTAn et al. (2022) 6.46 6.19 2.32 1.06 6.18
LLama3-8b-prompt 4.03 4.38 1.07 0.38 3.75
LLama3-8b-ICT 4.91 4.92 1.57 0.79 4.64
LLama3-70b-prompt 6.21 5.94 1.72 0.75 5.58
LLama3-70b-ICT 6.33 6.01 1.70 0.91 5.72
CMR 8.14 8.63 2.95 1.39 7.45
Table 1: The table shows the performance of different models on various automatic evaluation metrics. Our CMR model shows significant improvements over the baseline models. For the specific settings of the contrastive model, please refer to the appendix.
To evaluate the performance of our proposed CMR model, we compared it against several state-of-the-art models for dialogue response generation, including both general-purpose and multi-party dialogue models. Specifically, we considered BART Lewis et al. (2020) and T5 Raffel et al. (2020), both powerful transformer-based models known for their strong performance across NLP tasks, including dialogue generation. We also evaluated HeterMPC Gu et al. (2022a), which uses a heterogeneous graph neural network to capture complex interactions in multi-party dialogues, and MADNet Gu et al. (2023), which applies fully connected graph neural networks to address scenarios with missing addressee labels. Additionally, we included CONT An et al. (2022), which leverages contrastive learning to improve text generation by constructing contrastive examples from the model’s own predictions. Lastly, we used LLama3, a popular large language model with enhanced architecture and training techniques, as a baseline, testing various sizes and employing both prompt-based and in-context learning to assess its performance on our dataset.
4.3 Implementation Details
Models Metrics Rel. Coh. Eng. SA.
Bart 1.42 3.01 1.23 1.25
T5 1.84 2.67 1.66 1.23
HeterMPC 1.46 2.67 1.56 1.52
MADNet 2.12 2.56 2.13 1.87
CONT 2.61 3.21 3.04 2.27
CMR 3.06 3.40 3.09 2.88
Table 2: The table shows the performance of different models on various human evaluation metrics. Our CMR model shows significant improvements over the baseline models.
Our CMR model is implemented using PyTorch, with T5-base serving as the backbone. The learning rate is set to $4e^{-5}$ and we use the AdamW optimizer. The training procedure is divided into two stages. In Stage One, the encoder is trained with contrastive learning for 10 epochs with a batch size of 4. For each batch, 7 negative samples are used, with 4 samples coming from the same conversation and 3 from other samples within the same batch. After completing Stage One, Stage Two involves joint training with response generation and contrastive learning for 2 epochs, also with a batch size of 4. In this stage, 8 negative samples are used, distributed as follows: 1 sample from the previous utterance by the target speaker, 2 samples from other speakers within the same conversation, 3 samples from the same batch, and 2 model-generated beam search samples. The temperature $\tau$ is setting to 0.1. The model training was conducted on an RTX Titan GPU.
4.4 Automatic Evaluation Results
As shown in Table 1, to evaluate the performance of our proposed CMR model, we conducted experiments comparing it against several state-of-the-art models: BART, T5, CONT, HeterMPC, and MADNet. The evaluation metrics used include F1 Score, BLEU, and ROUGE-L, which provide a comprehensive assessment of the quality, relevance, and fluency of the generated responses. what is more, the results are averaged over five runs.
Compared to its backbone model, T5-base, CMR shows substantial improvements across all metrics, with some metrics nearly doubling. In particular, compared to CONT, which is also a contrastive learning-based response generation model, our model also shows significant improvements. Specifically, the F1 Score of CMR is 26.02% higher than that of CONT. For BLEU-1, BLEU-2, BLEU-3 and ROUGE-L, CMR achieves relative improvements of 39.26%, 27.16%, 31.13% and 20.58%, respectively. These results suggest that our approach to integrating contrastive learning with multi-party dialogue response generation is highly effective.
4.5 Human Evaluation Results
In addition to automatic evaluation, we conducted human evaluations (Table 2) to assess the quality of responses generated by our CMR model compared to baseline models. We used four metrics for evaluation: Relevance, Coherence, Engagement, and Speaker Appropriateness. Here, Speaker Appropriateness checks if the response matches the target speaker’s style and persona. Annotators rated each response on a scale of 1 to 5, with dialogue contexts specifically selected to require responses from one of the six main characters in "Friends." Annotators, blind to the model that generated each response and were instructed to reviewing training set utterances to familiarize themselves with speaking style of the main characters.
Models Metrics F1 BLEU-1 BLEU-2 BLEU-3 Rouge-L
CMR 8.14 8.63 2.95 1.39 7.45
-w/o SS1 7.37 7.64 2.55 1.26 6.77
-w/o SE1 7.83 7.99 2.61 1.31 7.02
-w/o Stage I 7.17 7.04 2.57 1.21 6.86
-w/o PU 7.88 7.89 2.74 1.19 7.09
-w/o SS2 6.93 6.96 2.48 1.07 6.41
-w/o SE2 7.05 6.89 2.68 1.34 6.64
-w/o BS 6.80 6.52 2.58 1.31 6.39
-w/o Stage II 6.41 6.32 2.24 1.04 5.91
Table 3: The table shows the performance of the CMR model and its variants with specific components removed.
As shown in Table 2, our model outperformed the baseline models in all metrics. Notably, general pre-trained models like BART and T5 often produced dull responses, with around 65% of their outputs resembling generic phrases like "I am sorry" or "yes yes yes." This resulted in lower scores for relevance, engagement, and speaker appropriateness, despite relatively high coherence scores due to their logical consistency. Meanwhile, graph-based models like HeterMPC and MADNet struggled without additional annotations, with HeterMPC frequently generating meaningless phrases across all samples. Finally, while graph structures helped improve relevance and coherence, particularly for HeterMPC and MADNet, contrastive learning in models like CMR proved more effective in capturing different speaking styles, demonstrating the unique advantages of each approach.
5 Analysis and Discussion
5.1 Ablation Study
To understand the contribution of each component of our CMR model, we conducted a series of ablation experiments. Each ablation removes a specific component or modifies the model’s training process, allowing us to observe the impact on performance. The results of these ablation studies are summarized in Table 3.
5.1.1 Impact of SS
The removal of SS1 (same conversation’s other speakers’ utterances as negative samples in Stage 1) and SS2 (same conversation’s other speakers’ utterances as negative samples in Stage 2) both resulted in performance drops across all metrics. Specifically, the F1 score decreased by 0.77 points when SS1 was removed and by 1.21 points when SS2 was removed. These results indicate the importance of incorporating context-specific negative samples to enhance the model’s ability to distinguish between relevant and irrelevant utterances.
5.1.2 Impact of Stage I
Removing the entire Stage I resulted in a significant performance drop across all metrics. The F1 score decreased by 0.97 points, BLEU-1 by 1.59 points, BLEU-2 by 0.38 points, BLEU-3 by 0.18 points, and ROUGE-L by 0.59 points. This demonstrates that the contrastive learning stage is crucial for improving the model’s understanding of dialogue context and speaker styles.
$N_{1},T_{1},N_{2},T_{2},P,B$ Metrics F1 BLEU-1 BLEU-2 BLEU-3 Rouge-L
$1,1,1,1,1,1$ 7.84 8.00 2.63 1.40 7.09
$2,2,2,1,1,1$ 7.95 8.47 2.69 1.37 7.20
$3,3,3,1,1,1$ 7.96 8.44 2.76 1.40 7.19
$\mathbf{4,4,4,2,1,2}$ 8.14 8.63 2.95 1.39 7.45
$5,5,5,3,1,3$ 8.17 8.64 2.87 1.41 7.51
$6,6,6,3,2,3$ 8.20 8.76 2.89 1.37 7.54
$8,8,8,4,3,4$ 8.02 8.58 2.77 1.34 7.20
Table 4: Performance of the CMR model with different numbers of negative samples ($N_{1}$, $T_{1}$, $N_{2}$, $T_{2}$, $P$, $B$). Increasing the number of negative samples initially improves performance, but beyond a certain point, the results start to fluctuate, indicating that excessively high numbers of negative samples can introduce irrelevant or low-quality negatives, which can degrade the model’s effectiveness.
5.1.3 Impact of PU and BS
In Stage 2, removing PU (previous utterance by the target speaker as a negative sample) and BS (beam search generated utterances as negative samples) led to noticeable declines in performance. These results suggest that PU helps the model achieve a fine-grained understanding of the dialogue context, while BS significantly enhances the overall quality of the generated responses.
5.1.4 Impact of Stage II
Removing the entire Stage II resulted in the most significant performance drop across all metrics. The F1 score decreased by 1.73 points, BLEU-1 by 2.31 points, BLEU-2 by 0.71 points, BLEU-3 by 0.35 points, and ROUGE-L by 1.54 points. This underscores the critical role of the joint training phase in fine-tuning the model’s response generation capabilities. Without Stage II, the model struggles to maintain the coherence and relevance of responses, demonstrating the necessity of this stage for achieving high-quality dialogue generation.
5.2 Impact of the Number of Negative Samples
The number of negative samples is crucial for contrastive learning. Therefore, we explored the impact of varying the number of negative samples on the performance of CMR. As shown in Table 4, increasing the number of negative samples initially improves the experimental results. However, beyond a certain point, the performance begins to fluctuate and does not continue to improve steadily. This fluctuation may be due to the inclusion of more irrelevant or less similar negative samples as the number of negatives increases. For instance, as the value of k increases in beam search-generated top-k negative samples, more less optimal negative samples are included, which can negatively impact the effectiveness of contrastive learning.
5.3 Visualization of the Impact of Speaker aware in Stage I
Figure 3: Attention weights of the $<$/s$>$ token in the encoder for (a) the CMR model with Stage I contrastive learning, and (b) the CMR model without Stage I contrastive learning. The attention weights are more focused on contextually relevant tokens such as names and pronouns in (a), indicating improved context understanding after Stage I.
To further understand the impact of Speaker aware in Stage I on the performance of the CMR model, we visualized the representations learned by the model with and without Stage I, as shown in Figure 3. Specifically, we visualized the attention weights of the $<$/s$>$ token to other tokens in the encoder. The reason for choosing the $<$/s$>$ token, as mentioned earlier, is based on our previous experiments. We found that during generation, the cross-attention weights of the decoder in the T5 model are primarily concentrated on the $<$/s$>$ token in the encoder. This indicates that the $<$/s$>$ token acts as an information aggregator in encoder, and its attention weights to other context tokens can effectively reflect the model’s primary focus. As shown in Figure 2, the attention weights of the $<$/s$>$ token in the CMR model with Stage I show a more concentrated pattern. The attention is focused on specific tokens that are contextually relevant, such as names (“Mona", “Ross") and pronouns (“I," “you"), indicating a better understanding of the dialogue context.
6 Conclusion
In this paper, we proposed CMR, a novel contrastive learning-based multi-party dialogue response generation model. Our approach leverages a two-stage training process to enhance the model’s understanding of dialogue context and speaker styles. In Stage One, we employ contrastive learning to differentiate between relevant and irrelevant utterances, significantly improving the model’s ability to comprehend complex multi-party dialogues. Stage Two combines response generation with contrastive learning, further refining the model’s capability to produce contextually relevant and coherent responses. In conclusion, our CMR model represents a significant advancement in multi-party dialogue response generation, offering a robust framework that effectively integrates contrastive learning to tackle the intricate dynamics of multi-party conversations. We believe that our findings will inspire further research and development in this challenging area, ultimately leading to more advanced and capable dialogue systems.
7 Limitations
While the CMR model demonstrates strong performance in multi-party dialogue response generation, several limitations must be acknowledged.
  • The use of contrastive learning, particularly in a multi-stage training process, results in significantly increased training times compared to simpler models. Constructing negative samples and distributed training further adds to the computational cost.
  • Although the model performs well on the Friends dataset, which has clear and well-defined speaker identities, its ability to generalize to other domains or dialogues with less structured or noisier data has not been thoroughly explored.
  • While the model effectively captures speaker-specific characteristics in datasets with distinct characters, such as Friends, it may not perform as well in scenarios where speakers frequently change or where speaker-specific features are less defined. This limits its effectiveness in more spontaneous, open-domain dialogues.
Appendix A The algorithm of entire training process
See Algorithm 1.
Algorithm 1 Training CMR Model
1:Stage One: Encoder Training
2:for epoch in num_epochs_stage_1 do
3:   for batch in dataloader do
4:      Initialize $loss_{stage\_1}\leftarrow 0$
5:      for dialogue, speaker in batch do
6:         Select query $q$ from dialogue
7:         Select positive sample from same speaker
8:         Select negative samples from other speakers and batch
9:         Compute InfoNCE loss
10:         $loss_{stage\_1}\leftarrow loss_{stage\_1}+loss_{InfoNCE}$
11:      end for
12:      Backpropagation and optimization
13:   end for
14:end for
15:Stage Two: Joint Training
16:for epoch in num_epochs_stage_2 do
17:   for batch in dataloader do
18:      Initialize $loss_{gen}\leftarrow 0$, $loss_{cl}\leftarrow 0$
19:      for dialogue D, target speaker $s_{t}$ in batch do
20:         Generate response $\hat{r}$ and negative samples from beam search
21:         Get gold response $r$
22:         Compute generation loss: $loss_{gen}\leftarrow loss_{gen}+\mathcal{L}_{gen}$
23:         Select query $q\leftarrow D$
24:         Select positive sample $r$
25:         Select negative samples from previous utterance, other speakers, batch, beam search
26:         Compute InfoNCE loss: $loss_{cl}\leftarrow loss_{cl}+\mathcal{L}_{cl}$
27:      end for
28:      Combine losses: $loss_{stage\_2}\leftarrow loss_{gen}+\lambda\cdot loss_{cl}$
29:      Backpropagation and optimization
30:   end for
31:end for
Appendix B Clean Friends Dataset
The Friends dataset we obtained included many low-quality samples. Some samples contained numerous meaningless and repetitive phrases, while others were too short and involved only two speakers. Additionally, some samples were fragmented and lacked context, making them unsuitable for training a robust dialogue model.
Bad Words
yes okay hey oh
no hello yeah well
aha ok i am
fine too bye is
good hm alright really
you are ohh uh
huh ha what go
hi thanks right many
and ohhh could can
that
Table 5: List of Bad Words
To address these issues, we first set a list of “bad words," as shown in Table 5. If a sample’s response contained these bad words in more than 30% of its content, the sample was removed. Additionally, we removed samples where the dialogue involved only two speakers. Finally, we manually removed samples that lacked sufficient context.
Appendix C Configuration of Comparison Models
The original HeterMPC and MADNet models use BART-base as their backbone model. We first tested their performance with this configuration. Then, we replaced their backbone model with BART-large. During this modification, we adjusted the “num_gat_heads" parameter in their configuration files to 8, as the original setting of 6 was not divisible by the 1024-dimensional hidden state of BART-large. The results reported in the paper use BART-large as the backbone because both models achieved better performance with this configuration.
For the configurations of LLaMA-3 using Prompt and in-context learning, refer to Table 6. In our experiments, the prompt shown in the table yielded the best results and is used in the reported experiments. The in-context learning configuration involved testing with 1 to 6 samples, and we found that using 3 samples provided the best performance.
Prompt $X_{t}$. You participated in
this conversation as $S_{t}$.
Please provide your
response.
[…]
Table 6: Configurations for LLaMA-3 using Prompt and In-context Learning. $t$ represents the target sample, and $i$ represents a random sample from the training set.